OpenStudy (anonymous):

The population of a European city increases by 2% annually. If the initial population was 50,000, what is the expected population in 10 years if this trend continues? A. 6,095 people B. 60,950 people C. 309,587 people D. 371,504 people

OpenStudy (tkhunny):

Trying again? Start: 50000 End of Year 1: 50000*1.02 = 51,000 End of Year 2: 50000*1.02^2 = 50000*1.0404 = 52,020 End of Year 3: 50000*1.02^3 = 50000*1.061208 = 53,060.4 <== This should make you a little uncomfortable. ... End of Year 10: You tell me.

OpenStudy (anonymous):

i got that by the end of the 10th year it would be 60,949.721 so im guessing the answer would be B.60,950 people

OpenStudy (tkhunny):

Why is that a guess? Go forth and tackle the next one.